# SPHINX: Structural Prediction using Hypergraph Inference Network

### v1

@@ 4.2 NBA Dataset
To evaluate our model on a real-world dataset, we use the NBA SportVU dataset, containing information about the movement of the players during the basketball matches. Each example contains $11$ trajectories, $5$ players from each team and one trajectory for the ball. Similar to the synthetic setup, the hypergraph predictor receives as node features the first part of each trajectory (the initial $5$ steps) and the goal is to predict the next $10$ steps. While our model is not especially designed as a tracking systems, the dynamics followed by the basketball game contains higher-order relations that are difficult to identify, thus representing a good testbed for our hypergraph inference method.
@@
Ablation study. We assess our core design choices on the real-world dataset as well. In Table 1 we are reporting the average displacement error (ADE) and final displacement error (FDE) metrics. Similar to the particle simulation datasets, we observe that predicting the hyperedges sequentially outperforms the simultaneous approach used by standard differentiable clustering algorithms. Moreover, the $k$-subset sampling prove to be clearly superior to the Gumbel-Softmax used in the previous methods.
@@
Comparison to recent methods. In Table 2 we compare against other structure-based methods in the literature including graph-based methods: NRI [19], DNRI [22], EvolveGraph [51], STGAT [52], Trajectron++ [53] and hypergraph-based methods: higher-order-only EvolveHypergraph, the full EvolveHypergraph model [43] and GroupNet [42]. Our method improves the short and medium-term prediction, while obtaining competitive results on the long-term prediction. Note that, compared to the other hypergraph-based methods, our model do not requires any auxiliary loss function or additional pairwise predictions.
@@
The metrics used in the recent literature ($\text{minADE}_{20}$ and $\text{minFDE}_{20}$) takes into account only the best trajectory from a pool of $20$ sampled trajectories. Recent work [54] shows that these metrics suffer from a series of limitations, favouring methods that produce diverse, but not necessarily accurate trajectories. Thus, in Table 1, we also compared against the top hypergraph-based method using the single-trajectory prediction metrics, ADE and FDE (the code for EvolveHypergraph is not publicly available). The results clearly show that for single-trajectory prediction our method obtains better results both on short and long-term prediction.
@@
Implementation details. In all our experiments we follow the train-valid-test split from [42]. For the observed trajectory, besides the coordinates of each player, the node features are enhanced with the velocity computed based on consecutive frames. For the unobserved trajectory, the velocity is computed based on the predicted position. The models are trained for 300 epochs, using Adam optimiser with learning rate 0.001, decreased by a factor of 10 when reaching a plateau. Similar to the synthetic setup, we are treating the algorithm used to perform $k$-sampling as a hyperparameter, experimenting with both AIMLE [9] and SIMPLE [8] algorithm. More details for the hyperparameters tuned in our experiments are presented in the Supplementary Material.
Table 2: Performance on the NBA dataset in terms of $\text{minADE}_{20}$ and $\text{minFDE}_{20}$ metrics. Our model performing higher-order processing using a learnable hypergraph structure prove to be beneficial for trajectory prediction tasks, obtaining competitive performance.
Time NRI DNRI EvolveGraph STGAT Trajectron++ EvolveHgraph EvolveHgraph GroupNet SPHINX
HO only full
1sec 0.51/0.74 0.59/0.70 0.35/0.48 0.45/0.66 0.44/0.67 0.49/0.74 0.33/0.49 0.34/0.48 0.30/0.43
2sec 0.96/1.65 0.93/1.52 0.66/0.97 0.87/1.41 0.79/1.18 0.95/1.68 0.63/0.95 0.62/0.95 0.59/0.94
3sec 1.42/2.50 1.38/2.21 1.15/1.86 1.28/2.08 1.51/2.49 1.44/2.27 0.93/1.36 0.87/1.31 0.88/1.38
4sec 1.86/3.26 1.78/2.81 1.64/2.64 1.69/2.66 2.09/3.52 1.91/3.08 1.21/1.74 1.13/1.69 1.16/1.74
@@ 5 Conclusion
In this paper we are introducing Structural Prediction using Hypergraph Inference Network (SPHINX), a model for unsupervised hypergraph inference that enables higher-order processing in tasks where hypergraph structure is not provided. The model adopts a global processing in the form of sequential soft clustering for predicting hyperedge probability, paired with a $k$-subset sampling algorithm for discretizing the structure such that it can be used with any hypergraph neural network decoder and eliminating the need for heavy optimisation tricks. Overall, the resulting method is a general and broadly applicable solution for relational processing, facilitating seamless integration into any real-world systems perceived to necessitate higher-order analysis.
@@ Acknowledgment
The authors would like to thank Charlotte Magister, Daniel McFadyen, Alex Norcliffe and Paul Scherer for fruitful discussions and constructive suggestions during the development of the paper. Iulia Duta is a PhD student funded by a Twitter scholarship.
Appendix: SPHINX Structural Prediction using Hypergraph Inference Network
This appendix contains details related to our proposed Structural Prediction using Hypergraph Inference Network (SPHINX) model, including broader impact and potential limitations, qualitative visualisations from our model and the baselines we compared against in this paper, details about the proposed dataset, information about the implementation of the model and additional experiments referred in the main paper.
@@
  • Section B contains visualisation of the discovered hypergraph structure and the predicted trajectory for our proposed model and different variations of it.
  • Section C provides more details about the proposed synthetic dataset, further information about the method and its training process.
  • Section D presents additional ablation studies showing the influence of varying the number of hyperedges in our model, but also a more detailed analysis of the synthetic experiments presented in the paper, including the numerical results.
@@ Appendix A Broader Impact & Limitations
In this paper, we are proposing a framework for processing higher-order interactions in scenarios where a ground-truth hypergraph connectivity is either too expensive or not accessible at all. We are showing that our model is capable of recovering the latent higher-order structure without supervision at the hypergraph level, thus improving the final, higher-level performance. While we mainly tested on trajectory prediction tasks, our model is general and versatile and can be applied to any scenarios where we suspect capturing and modeling higher-order relations can be beneficial. Consequently, we believe that our model doesn’t have any direct negative societal impact. Moreover, the method has the advantage of allowing us to inspect the learned latent hypergraph structure (as we can see in Section B of this Supplementary Material). This offer an advantage in terms of interpretability, allowing us to better understand and address potential mistakes in the model’s prediction.
@@
In order to better align with the current literature, we mainly validate our model on the trajectory prediction task. However, in the current form, our model predicts a single hypergraph structure for the entire trajectory. This is well suited for the synthetic benchmark, where the dynamics for each example is determined by a single hypergraph structure, that does not change in time. On the other hand, in real-world setups, such as the NBA dataset, the higher-order interactions between the players might change along the game. Thus, having a distinct, dynamic hypergraph structure, that evolves in time, could be beneficial. However, solving trajectory prediction is not the main goal for our model. Instead, our purpose is to have a general model, that allows us to infer and process higher-order interaction. We believe that adapting SPHINX to more specific scenarios, such as creating a dynamic, evolving structure is an interesting area for future work.
Appendix B Visualisations
@@
One of the advantages brought by our method is providing a discrete latent hypergraph structure, that can be easily visualise, offering a higher level of interpretability compared to attention-based methods. Figure 6 shows the trajectories and latent structure learned by our model on some examples from the Two-Triangles and NBA datasets. We are visualising the ground-truth trajectories, the trajectories and structure predicted by our full model, and the trajectories and learned structure corresponding to the model used in our ablation study: SPHINX w/o sequential and SPHINX w Gumbel.
@@
Appendix C Experimental details
C.1 Datasets details
Figure 5: Examples of $25$-steps trajectories from the Particle Simulation datasets for both the One-Triangle and the Two-Triangles variants. The highlighted triangles represent the 3-order interactions used to generate the trajectories.
@@ C.1 Datasets details
We create two variants of the dataset. A simpler one, containing a single higher-order interaction per trajectory (denoted as One-Triangle), and a more challenging one containing two higher-order interaction per trajectory (denoted as Two-Triangles). Each dataset contains $1000$ training, $1000$ validation and $1000$ test trajectories.
Some examples of the datasets are depicted in Figure 5.
Figure 6: Visualisation of the trajectories predicted by our full model (SPHINX) and the two variations used in the ablation study (SPHINX w/o sequential and SPHINX w Gumbel) on the Two-Triangles and NBA dataset. In the ground-truth column the highlighted polygons represent the true connectivity, while for the models they represent the discovered hyperedges. Both using Gumbel-Softmax and dropping the sequential prediction clearly impact the predicted hypergraphs. On the other hand, our model manage to discover diverse structures, close to the ground0truth ones in the synthetic scenario.
C.2 Implementation details
@@
The code for loading the NBA dataset is based on the official code for  [42] [^1] https://github.com/MediaBrain-SJTU/GroupNet - under MIT License. The code for the various $k$-subset sampling algorithms is based on the code associated with AIMLE [^2] https://github.com/EdinburghNLP/torch-adaptive-imle - under MIT License [9], IMLE [^3] https://github.com/uclnlp/torch-imle under MIT License [46] and SIMPLE [^4] https://github.com/UCLA-StarAI/SIMPLE - custom License, open for research [8] methods.
@@
Computational complexity. We analyse the computational complexity of the hypergraph-inference component for the SPHINX architecture (hypergraph predictor + sampling) when using the SIMPLE $k$-subset sampling algorithm. The overall complexity would be obtained by adding the complexity specific to the hypergraph neural network used to process the higher-order structure. If $N$ is the number of nodes, $K$ is the cardinality of the predicted hyperedges and $M$ is the number of inferred hyperedges, the vectorized complexity for the sampling module is $O(logN\times logK)$ and the complexity of slot-attention with $Q$ iterations and one slot is $O(Q\times N)$. Then, the overall complexity for predicting $M$ hyperedges is $O(M\times Q\times N+M\times logN\times logK))$.
Appendix D Additional experiments
Table 3: Ablation study on the One-Triangle dataset. Taking into account the higher-order interaction is crucial for predicting the trajectory. Both predicting the hyperedges sequentially and the constrained $k$-subset sampling are important for accurate hypergraph prediction.
Model 1 step 5 steps 25 steps
Golden standard 1.0e-8 1.0e-7 1.0e-6
Random structure 1.1e-4 1.1e-3 2.4e-2
No structure 1.1e-4 1.2e-3 2.4e-2
LSTM 2.0e-5 1.0e-1 1.7e-1
SPHINX w Gumbel 1.0e-4 1.0e-3 2.3e-2
SPHINX 3.6e-6 3.9e-5 1.1e-3
Table 4: Ablation study on the Two-Triangles dataset, reinforcing the conclusions from the One-Triangle dataset. Having a learnable hypergraph structure improves the trajectory prediction, while the constrained $k$-subset sampling is superior to Gumbel-Softmax. Moreover, the model further benefits from a sequential prediction of slots.
Model 1 step 5 steps 25 steps
Golden standard 4.0e-8 5.0e-7 9.0e-6
Random structure 1.1e-4 1.1e-3 2.0e-2
No structure 1.1e-4 1.2e-3 2.5e-2
LSTM 3.0e-5 9.5e-2 1.7e-1
SPHINX w/o sequen. 5.0e-5 5.0e-4 1.0e-2
SPHINX w Gumbel 1.0e-4 1.1e-3 2.5e-2
SPHINX 2.0e-5 2.0e-4 6.0e-3
D.1 Full results on Particle Simulation
Due to page limits, in the main paper we only report the bar-plot for our synthetic experiments. Here we also include the numerical values (see Table 3 and Table 4), together with two additional models: one LSTM baseline, similar to the one proposed in  [19], and a ground-truth-based oracle, to better understand the lower bound for our model.

### v2

@@ 4.2 Inductive setup
To evaluate our model on an inductive real-world dataset, where we need to predict a distinct hypergraph for each example, we use the NBA SportVU dataset, containing information about the movement of the players during basketball matches. Each example contains $11$ trajectories, $5$ players from each team and one trajectory for the ball. The hypergraph predictor receives as node features the first $5$ timesteps of each trajectory and the goal is to predict the next $10$ steps. While our model is not especially designed as a tracking system, the dynamics followed by the basketball game contain higher-order relations that are difficult to identify, thus representing a good testbed for our method.
@@
Ablation study. We assess our core design choices on the real-world dataset as well. In Table 1 we are reporting the average displacement error (ADE) and final displacement error (FDE) metrics. Similar to the particle simulation datasets, we observe that predicting the hyperedges sequentially outperforms the simultaneous approach (SPHINX w/o sequential model) used by standard differentiable clustering algorithms. Moreover, the $k$-subset sampling proved to be clearly superior to the Gumbel-Softmax used in the previous methods (SPHINX w gumbel model).
@@
Comparison to recent methods. In Table 3 we compare against other structure-based methods in the literature including graph-based methods: NRI (Fetaya et al., 2018), DNRI (Graber & Schwing, 2020), EvolveGraph (Li et al., 2020), STGAT (Huang et al., 2019), Trajectron++ (Salzmann et al., 2020) and hypergraph-based methods: higher-order-only EvolveHypergraph (H), the full EvolveHypergraph (H+G) (Li et al., 2022), GroupNet (Xu et al., 2022) and TDHNN (Zhou et al., 2023). Our method improves the short and medium-term prediction, while obtaining competitive results on the long-term prediction. Moreover, it is the only hypergraph-based method that does not require any auxiliary loss function or additional pairwise predictions.
@@
The metrics used in the recent literature ($\text{minADE}_{20}$ and $\text{minFDE}_{20}$) take into account only the best trajectory from a pool of $20$ sampled trajectories. Recently  (Greer et al., 2020) show that these metrics suffer from a series of limitations, favouring methods that produce diverse, but less accurate trajectories. Thus, in Table 1, we also compared against the existing hypergraph-based methods using the single-trajectory prediction metrics (ADE/FDE) (the code for EvolveHypergraph is not publicly available). The results clearly show that for single-trajectory prediction our method obtains better results on both short and long-term prediction.
@@
Implementation details. In all experiments we use the split from (Xu et al., 2022). For the observed trajectory the features are enhanced with the velocity. For the unobserved trajectory, the velocity is computed based on the predicted position. The models are trained for 300 epochs, using Adam with lr. 0.001, decreased by a factor of 10 when reaching a plateau. We treat the sampling algorithm as a hyperparameter, experimenting with AIMLE and SIMPLE.
4.3 Transductive setup
Most of the hypergraph inference methods are constrained to only predict a single hypergraph, which makes them unsuitable for the inductive setup. To allow direct comparison with all methods, we perform experiments on two transductive datasets for object classification. ModelNet40 (Wu et al., 2014) contains $12311$ objects of $40$ types, while NTU (Chen et al., 2003) contains $2012$ objects with $67$ types.
Comparison to recent methods. We compare against both static methods (where the structure does not adapt to the task) GCN, GAT (Veličković et al., 2018), HGNN, HGNN+ (Gao et al., 2023), HNHN (Dong et al., 2020), HyperGCN (Yadati et al., 2019) and dynamic hyperaph methods (where the structure is learned) DHGNN (Jiang et al., 2019), DeepHGSL (Zizhao et al., 2022), HSL (Cai et al., 2022), TDHNN. Table 2 shows that SPHINX learns a better latent structure, improving the final performance.
Note that, upon retraining the TDHNN model, we improved their official results. For fairness we report the improved performance (TDHNN*). However, qualitative evaluation showed that in all cases each hyperedge learned by TDHNN ends up containing all the nodes in the dataset. While the model still obtains a good downstream performance, it seems that this is the results of learning good attention coefficients rather than learning a useful sparse structure.
Implementation details. For all transductive experiments we adopt the split from  (Zhou et al., 2023) and train the model with three types of visual features: GVCNN (Feng et al., 2018), MVCNN (Su et al., 2015) or combining both of them. We are treating the number of slots and the cardinality of slots as hyperparameters. Unless otherwise stated, we use AllDeepSet model as hypergraph processor.
@@ 5 Conclusion
We are introducing SPHINX, a model for unsupervised hypergraph inference that enables higher-order processing in tasks where hypergraph structure is not provided. The model adopts a global processing in the form of sequential soft clustering for predicting hyperedge probability, paired with a $k$-subset sampling algorithm for discretizing the structure such that it can be used with any hypergraph neural network decoder and eliminating the need for heavy optimisation tricks. Overall, the resulting method is a general and broadly applicable solution for relational processing, facilitating seamless integration into various real-world systems perceived to necessitate higher-order analysis.
@@ Acknowledgment
The authors would like to thank Charlotte Magister, Daniel McFadyen, Alex Norcliffe and Paul Scherer for fruitful discussions and constructive suggestions during the development of the paper.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix
This appendix contains details related to our proposed Structural Prediction using Hypergraph Inference Network (SPHINX) model, including broader impact and potential limitations, qualitative visualisations from our model and the baselines we compared against in this paper, details about the proposed dataset, information about the implementation of the model and additional experiments referred in the main paper. We will soon release the full code associated with the paper.
@@
  • Section B presents an overview of the existing methods for hypergraph prediction, with a focus on a few qualities that we believe a general methods should fulfill.
  • Section C present the computational complexity of our model.
  • Section D contains visualisation of the discovered hypergraph structure and the predicted trajectory for our proposed model and different variations of it.
  • Section E provides more details about the proposed synthetic dataset, further information about the method and its training process.
  • Section F presents additional ablation studies showing the influence of varying the number of hyperedges in our model, but also a more detailed analysis of the synthetic experiments presented in the paper.
@@ Appendix A Broader Impact & Limitations
In this paper, we are proposing a framework for processing higher-order interactions in scenarios where a ground-truth hypergraph connectivity is either too expensive or not accessible at all. We are showing that our model is capable of recovering the latent higher-order structure without supervision at the hypergraph level, thus improving the final, higher-level performance. While we mainly tested on trajectory prediction tasks, our model is general and versatile and can be applied to any scenarios where we suspect capturing and modeling higher-order relations can be beneficial. Consequently, we believe that our model doesn’t have any direct negative societal impact. Moreover, the method has the advantage of allowing us to inspect the learned latent hypergraph structure (as we can see in Section D of this Supplementary Material). This offer an advantage in terms of interpretability, allowing us to better understand and address potential mistakes in the model’s prediction.
@@
On the inductive setup, in order to better align with the current literature, we mainly validate our model on the trajectory prediction task. However, in the current form, our model predicts a single hypergraph structure for the entire trajectory. This is well suited for the synthetic benchmark, where the dynamics for each example is determined by a single hypergraph structure, that does not change in time. On the other hand, in real-world setups, such as the NBA dataset, the higher-order interactions between the players might change along the game. Thus, having a distinct, dynamic hypergraph structure, that evolves in time, could be beneficial. However, solving trajectory prediction is not the main goal for our model. Instead, our purpose is to have a general model, that allows us to infer and process higher-order interaction. We believe that adapting SPHINX to more specific scenarios, such as creating a dynamic, evolving structure is an interesting area for future work.
Appendix B Overview of the existing dynamic hypergraph predictors
Table 4: Overview of the existing methods for hypergraph prediction.
Method inductive transductive differentiable discrete regularisation-free
[…]
HERALD (Zhang et al., 2022) requires initial hgraphs
Ada-Hyper (Yu et al., 2012)
DyHL (Zhang et al., 2018)
DHGNN (Jiang et al., 2019)
(D)HSL (Wang et al., 2024; Cai et al., 2022)
[…]
EvolveHypergraph (Li et al., 2022)
SPHINX (Ours)
For a better understanding of the field, we are offering an overview of the existing methods for dynamic hypergraph prediction, focusing on the three desiderata that we highlight in the introduction: 1) The model should be applicable on a broad types of tasks. For a general hypergraph prediction method, it should be adaptable for both the inductive and the transductive setup. 2) The model should be compatible with any hypergraph processing architecture. This requires the model to be differentiable and to produce a discrete output such that it can be easily plugged into any hypergraph decoder. 3) A powerful model should be easy to optimise. Most of the models requires additional regularisation terms or residual branches in order to optimise the hypergraph predictor well. Ideally, the model should allow to incorporate the hypergraph predictor component without negatively impacting the optimisation process.
For completeness, we are including also methods such as DeepHGSL, MHSL and HERALD, but these methods does not directly generate a hypergraph structure, but rather rewire an existing one. As we can see in Table 4, half of the models are unapplicable in the inductive setup mainly due to the fact that they are treating the structure as a parameter. The models marked with the symbol $\sim$ are only partially differentiable. Models such as top-k or threasholding, when using to infer a discrete structure, loses the gradient. Moreover, the table also suggests that most of the methods are having issue in optimising a sparse, discrete structure without incorporating various regularisation constraints.
The results presented in our main paper confirm the fact that SPHINX is a general and easy to adapt method, as it obtains good performance across various setups, without any additional regularisation.
Appendix C Computational complexity.
We analyse the computational complexity of the hypergraph-inference component for the SPHINX architecture (hypergraph predictor + sampling) when using the SIMPLE $k$-subset sampling algorithm. The overall complexity would be obtained by adding the complexity specific to the hypergraph neural network used to process the higher-order structure. If $N$ is the number of nodes, $K$ is the cardinality of the predicted hyperedges and $M$ is the number of inferred hyperedges, the vectorized complexity for the sampling module is $O(logN\times logK)$ and the complexity of slot-attention with $Q$ iterations and one slot is $O(Q\times N)$. Then, the overall complexity for predicting $M$ hyperedges is $O(M\times Q\times N+M\times logN\times logK))$.
Appendix D Visualisations
@@
One of the advantages brought by our method is providing a discrete latent hypergraph structure, that can be easily visualise, offering a higher level of interpretability compared to attention-based methods. Figure 6 shows the trajectories and latent structure learned by our model on some examples from the Particle Simulation and NBA datasets. We are visualising the ground-truth trajectories, the trajectories and structure predicted by our full model, and the trajectories and learned structure corresponding to the model used in our ablation study: SPHINX w/o sequential and SPHINX w Gumbel.
@@
Insights on the inferred hypergraph.
Evaluating the accuracy of predicted latent hypergraphs in real-world data is a critical but challenging problem, that remains an open problem. Our datasets lack annotations for higher-order structures, and even if they would have existed, there’s no guarantee they would be optimal for the task.
However, to understand the type of hypergraph predicted by our model in the real-world datasets, we analyze the NBA dataset, an inductive dataset that is easier to visualize and interpret. For a model with 6 hyperedges and 4 nodes per hyperedge:
  • the average node degree for the predicted hypergraphs are $[1.97,1.98,1.99,1.99,1.98,1.95,1.96,1.96,1.96,1.95,4.29]$. The first $10$ nodes correspond to players- model focuses uniformly on all of them; while the last node represents the ball- the model learns to include the ball in most hyperedges. Thus the predicted structures are star-like hypergraphs, aligning with our intuition that player movement is highly influenced by the ball’s position.
  • the average percentage of duplicated hyperedges across examples is $2.74\%$.
This shows a high diversity in our predictions, confirming that the model captures not just the most likely connections but also a scene-specific structure.
Appendix E Experimental details
E.1 Datasets details
Figure 4: Examples of $25$-steps trajectories from the Particle Simulation datasets. The highlighted triangles represent the 3-order interactions used to generate the trajectories.
@@ E.1 Datasets details
The dataset contains two higher-order interaction per trajectory. The dataset split constitutes of $1000$ trajectories for training, $1000$ forvalidation and $1000$ for test.
Some examples of the dataset are depicted in Figure 4.
E.2 Implementation details
@@
Figure 5: Experiments on Particle Simulation dataset investigating the importance of choosing the correct number of predicted hyperedges. Having less hyperedges than the real number (two) negatively impact the performance. However, when increasing the number of hyperedges above the true value, the performance does not deteriorate, indicating that the model is robust at discovering the hypergraph structure, as long as enough slots are provided.
@@
The code for loading the NBA dataset is based on the official code for  (Xu et al., 2022) [^4] https://github.com/MediaBrain-SJTU/GroupNet - under MIT License. The code for the various $k$-subset sampling algorithms is based on the code associated with AIMLE [^5] https://github.com/EdinburghNLP/torch-adaptive-imle - under MIT License (Minervini et al., 2023), IMLE [^6] https://github.com/uclnlp/torch-imle under MIT License (Niepert et al., 2021) and SIMPLE [^7] https://github.com/UCLA-StarAI/SIMPLE - custom License, open for research (Ahmed et al., 2023) methods.
@@
Figure 6: Visualisation of the trajectories predicted by our full model (SPHINX) and the two variations used in the ablation study (SPHINX w/o sequential and SPHINX w Gumbel) on the Particle Simulation and NBA dataset. In the ground-truth column the highlighted polygons represent the true connectivity, while for the models they represent the discovered hyperedges. Both using Gumbel-Softmax and dropping the sequential prediction clearly impact the predicted hypergraphs. Gumbel-Softmax struggles to produce a sparse structure, while the lack of sequentiality leads to both predicted hyperedges containing the same set of nodes. On the other hand, our model manage to discover diverse structures, close to the ground-truth ones in the synthetic scenario.
Appendix F Additional experiments
F.1 Description of the models used in Particle Simulation